OpenStudy (anonymous):

sec^4theta-8sec^2(theta)*tan(theta)+16tan^2(theta)=0 if theta=?

myininaya (myininaya):

\[\sec^2(\theta)-8\sec^2(\theta)\tan(\theta)+16\tan^2(\theta)=0\] let's just write in terms of \[\tan(\theta)\] first of all recall that \[\sin^2(\theta)+\cos^2(\theta)=1\] dividing both sides by \[\cos^2(\theta)\] will give us the identity we need \[\tan^2(\theta)+1=\sec^2(\theta)\] so now we can write our equation as \[(\tan^2(\theta)+1)^2-8(\tan^2(\theta)+1)\tan(\theta)+16\tan^2(\theta)=0\] now lets multiply a little \[\tan^4(\theta)+2\tan^2(\theta)+1-8\tan^3(\theta)-8\tan(\theta)+16\tan^2(\theta)=0\] if any combine like terms \[\tan^4(\theta)-8\tan^3(\theta)+18\tan^2(\theta)-8\tan(\theta)+1=0\] now lets use a substitution to make this easier on our eyes let \[u=\tan(\theta)\] so we have \[u^4-8u^3+18u^2-8u+1=0\] now we need to find the zeros of this polynomial \[1-8+18-8+1=-7+18-7=-14+18=4 \neq 0\] \[1+8+18+8+1 \neq 0\] so we have that 1 or -1 don't work so that means we have irrational solutions for this polynomial

OpenStudy (anonymous):

its ans is \[\Theta=5\pi/12 \Theta=\Pi /12, \]

OpenStudy (anonymous):

let t=tan(theta), the given equation became t^4-8t^3+18t^2+1=0 x=t+1/t =)x^2-8x+16=0 x=4 t=2+-\[\sqrt{3}\]

OpenStudy (anonymous):

its very simple problem to solv

myininaya (myininaya):

wait our polynomials are different i have an extra term let me look at this

myininaya (myininaya):

\[u^4-8u^3+18u^2-8u+1=0 \] i see this let me know if i made a mistake i know in my first line i typed it wrong but i fixed it when i was working on it above